OpenStudy (anonymous):

On a trip, the fare was 50¢ for each adult and 25¢ for each child. If 30 passengers paid $12.25, how many adults and children went? Is this the right way to set up the equation? 50x + 25(30 - x) = 1225

OpenStudy (anonymous):

it is, yes..

OpenStudy (anonymous):

D+C=30 C= 30-D 12.25=.5D+.25C 12.25=.5D+.25(30-D)

OpenStudy (anonymous):

I'm not sure lol

OpenStudy (anonymous):

Not quite.... 25x + 50x = 30(1225) is the answer.